# Periodic robust robotic rock chop via virtual model control

### v1

@@ VI-A Forces, velocities, and frequency of successful cut
We evaluated the versatility of the controller by cutting a variety of vegetables (spring onion, cucumber, carrot, courgette, and potato) under two conditions: (i) using a single set of control parameters for all food (uniform parameters), and (ii) with parameters tuned for each food type (food-tuned parameters), with the goal of minimizing the applied force. In fact, reducing the cutting force helps to achieve precise cuts and prolongs the lifespan of both the knife and the cutting board. Larger items, such as cucumber, courgette, and potato, were cut in half to make them suitable for rocking and ensure stable placement on the cutting board. All vegetables were securely held to prevent motion during cutting. Results are summarized in Fig. 10.
To quantitatively evaluate the cutting process, we have used the following metrics. Let $\mathcal{C}_{\mathrm{cut},i}=\{t_{1},\dots,t_{N_{i}}\}$ denote the times at which a sample measure is taken during the cutting phase $i$. These belong to the time interval in which $\boldsymbol{p}_{b}$ is connected to the reference $\boldsymbol{r}_{2,1}$, as shown in Fig. 3(c) (left). The average and peak forces during the $i$th cycle are computed from the force-torque sensor reading, mounted at the robot wrist
@@
When using a single set of control parameters for all foods, the force profiles in the $z$-axis (Fig. 10(a)) were similar across all food types despite significant differences in stiffness. This is because the control parameters were chosen to apply a sufficiently high force so that the intrinsic properties of the food had minimal impact on the resulting behavior. Peak force is typically observed near the transition from cut to raise and varies with the energy absorption characteristics of each vegetable. The level of ‘residual’ energy after cut typically results in a higher contact force before the knife is retracted. With parameters tuned specifically for each food type, the force required for cutting spring onion and cucumber is noticeably reduced, while high force levels are needed for other vegetables. Fig. 1 shows the sliced vegetables at the end of experiments. By visual inspection, both (i) uniform and (ii) food-specific parameter selection produce no noticeable difference in slice quality or accuracy.
@@
VI-B Cut accuracy
@@
VI-C Cut robustness
The VMC approach demonstrates inherent robustness in handling environmental variations. As shown in Section VI-A, the same set of control parameters enabled successful cut of a variety of food items, from delicate spring onions to rigid potatoes. In addition to these results, we tested the controller using three different knives, as shown in Fig. 12(a). Knife 1 features an optimal blade curvature for rocking and has been used in all experiments so far. Knife 2 has a different blade geometry, which is still suitable for rocking. In contrast, knife 3 is a carving knife and should not be used for rocking. Results are shown in Fig. 12. To assess robustness, we consider a uniform set of parameters for all knives.
Although all knives exhibited rocking motion under identical control parameters, the differences in blade geometry affected energy dissipation and resulted in different rocking features. Knife 2 had a smaller blade than knife 1, resulting in a reduced energy dissipation through contact with the cutting board. This led to a higher residual energy after cutting, which turned into speed, that is, faster cuts. Similar considerations apply to knife 3. The progressive increase in force peaks for knives 2 and 3, respectively, can be explained in a similar way: knives 2 and 3 are moving faster than knife 1, leading to stronger forces at contact with the cutting board.
We also demonstrate the ability of the controller to handle different heights of the cutting board. Fig. 13(a) illustrates the experimental setup, where the nominal cutting board height was modified by adding MDF layers underneath. Fig. 13 shows that the rocking motion remained stable across all heights, without requiring any adjustments to the control parameters. The gradual decrease in peak force with increasing height can be explained by the knife having less distance to accelerate before switching.
Together, Figs. 12 and 13 provide a neat illustration of the difference between our virtual model controller and trajectory-based methods. The latter would require a precise knowledge of knife geometry and new path planning computations to take into account variations in cutting board heights. In contrast, because the rocking motion established by the virtual model controller is not programmed, it can easily comply with various settings.
(a) Three different knives.
(b) Knives 1, 2, and 3 rocking motions.
(c) Forces, velocities, and cut frequencies of the three knives.
Figure 12: Rocking motion with three different knives while using the same virtual mechanism and control parameters.
@@ VI-C Cut robustness
(b) Forces and velocities for different heights of the board.
Figure 13: Rocking with different cutting boards.
VI-D Platform independence
The proposed virtual mechanism is platform independent. We demonstrate this by implementing the same controller on a humanoid robot from RT Corp, Sciurus17. Unlike the Franka platform, Sciurus17 has a lower payload capacity and slower communication speed. This can be managed by re-tuning the parameters of the controller to ensure smaller forces and slower motions. Fig. 14 shows that the inherent limitations of the robotic platform led to reduced performance. However, successful rocking motion is achieved.
@@ VI-D Platform independence
VII Conclusion and future works
We have proposed a new virtual model controller for rock cutting. The rocking motion has been discussed theoretically, through energy considerations, and experimentally. Performance has been illustrated by cutting several vegetables, at several thicknesses and cut frequencies. Robustness has been illustrated by using three different knives and several cutting board heights. Notably, the controller does not require any trajectory planning and relies on minimal information, both at design and deployment stage, and delivers consistent and reliable cutting performance. Experiment videos are available online: https://youtu.be/67v-U4gp9_k.
Although these results clarify the flexibility and reliability of our control strategy, there are still significant directions for improvement. At level of integration, the most pressing one is handling the full cutting pipeline. This entails autonomous pick and place of the food, its orientation on the cutting board, and the capability of holding food in place during cutting. It seems also crucial for the robot to recognize different foods and knives, and to measure in real-time the result of its action, during and after cutting. More work is also needed at the level of the food-knife interaction. A key future direction is the systematic exploration of adaptation/learning of the control parameters, to improve performance. The idea is to exploit additional sensing, such as vision and tactile information, to optimize control parameters in real time.
Thanks to the intuitive design of our virtual mechanism, automatic control parameter adjustment can be performed at different levels of complexity. Here, for completeness, we discuss an initial study in which the spring stiffness $k_{2}$ is adjusted to achieve successful food separation while moderating the force exerted on food and board. The overall strategy is extremely simple. We increase $k_{2}$ if the knife does not achieve separation of food (by position sensing through robot and knife geometry). We reduce $k_{2}$ if we detect excess overshoot on the knife handle (i.e. if the knife handle moves past the critical reference). This leads to an error signal, $e$, capturing the mismatch between desired and achieved vertical position of the knife. We update $k_{2}$ every six cutting cycles, using the adaptation law
@@ VII Conclusion and future works
where $\alpha$ is the adaptation gain.
Figure 15 provides a brief illustration. All control parameters are set as in the A-2 Fast cut in Table I, except for $k_{2}$, which is dynamically adapted. $k_{2}$ starts from a low value, which is sufficient initially to cut through a carrot. From time $t=20$s, the carrot presents a harder consistency which requires adaptation to cut through as shown in Figure 15(b). $k_{2}$ increases and, after $t=35$s, adaptation is completed and cutting is reestablished. This brief illustration shows the potential of continuous learning combined with virtual model control. Here we have focused on a simple single parameter for simplicity, but we plan to develop a systematic approach in the future, leveraging vision and tactile sensing.
@@
TABLE I: Control parameters for each experiment. The following parameters remain the same across all experiments: $m_{a}=0.1$, $m_{\text{ori}}=0.1$.
(a) Performance
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
A-1 Uniform parameters 25 10 1 150 25 2 1200 50 10 0.43, -0.1 0.58, 0.4 0.02 0.15 1 30 17
A-1 Spring onion 10 10 1 65 25 2 1200 50 10 0.43, -0.1 0.58, 0.4 0.02 0.2 0.01 30 1.7
A-1 Cucumber 10 10 1 75 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Courgette 10 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Carrot 20 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
A-1 Potato 20 10 1 105 25 2 1200 50 10 0.43, -0.1 0.83, 0.5 0.02 0.25 0.01 30 1.7
(b) Performance
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
A-2 Fast cut (0.94 cut/s) 45 10 1 200 30 2 800 50 10 0.43, -0.1 0.73, 0.35 0.035 0.145 0.01 250 4
A-2 Slow cut (0.51 cut/s) 25 10 1 120 25 2 800 50 10 0.43, -0.1 0.58, 0.4 0.02 0.15 1 30 17
(c) Robustness
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
B Robustness 30 10 1 130 25 2 800 50 10 0.43, -0.1 0.63, 0.4 0.01 0.2 0.005 30 3
(d) Platform independence, Sciurus
Exp $k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $m_{b}$ $k_{b}$ $c_{b}$
C Sciurus 10 2 0.3 20 10 0.1 30 10 1 0.4, -0.1 0.6, 0.6 0.047 0.38 0.8 110 50

### v2

@@ 6.1 Forces and velocities
To evaluate the controller’s versatility, we conducted cutting experiments on a variety of vegetables (spring onion, cucumber, carrot, courgette, and potato) under two conditions: (i) a uniform parameter set applied to all food types, and (ii) food-tuned parameters manually optimized to minimize contact force. Reducing the cutting force helps to achieve precise cuts and prolongs the lifespan of both the knife and the cutting board. Larger items, such as cucumber, courgette, and potato, were cut in half to make them suitable for rocking and ensure stable placement on the cutting board. All vegetables were securely held to prevent motion during cutting. Results are summarized in Fig. 8. The rocking motion was initially validated through simulation but parameter tuning was performed on the physical robot to account for the complexities of modeling contact forces during cutting. Although we explore the influence of these parameters on performance, a systematic optimization study remains beyond the scope of this work. We refer the reader to Larby and Forni (2025) for an optimization framework tailored to virtual model control.
To quantitatively evaluate the cutting process, we have used the following metrics. Let $\mathcal{C}_{\mathrm{cut},i}=\{t_{1},\dots,t_{N_{i}}\}$ denote the set of sampling times during the cutting phase $i$. These belong to the time intervals in which $\boldsymbol{p}_{b}$ is connected to the reference $\boldsymbol{r}_{2,1}$, as shown in Fig. 3(c) (left). The average and peak forces during the $i$th cycle are computed from the force-torque sensor reading, mounted at the robot wrist
@@
When using a single set of control parameters for all foods, the force profiles in the $z$-axis (Fig. 8(a)) were similar across all food types despite significant differences in the material properties of different food. This is because the control parameters were chosen to apply a sufficiently high force so that the intrinsic properties of the food had minimal impact on the resulting behavior. Peak force is typically observed near the transition from cut to raise and varies with the energy absorption characteristics of each vegetable. Higher levels of ‘residual’ energy after cut typically results in a higher contact forces before the knife is retracted. With parameters tuned specifically for each food type, the force required for cutting spring onion and cucumber is noticeably reduced, while high force levels are needed for other vegetables. In terms of tuning strategy, stiffness of springs was uniformly reduced (increased) to reduce (increase) contact forces. The mass–spring–damper parameters ($m_{b},k_{b},c_{b}$) were adjusted accordingly to preserve cutting frequency. For smaller items, the knife’s lifting height was reduced by increasing the switching tolerance $\delta_{2}$, maintaining comparable cutting frequencies even with lower cutting velocities. For larger items, the reference position for raising the knife, $\boldsymbol{r}_{2,2}$, was set higher to accelerate the upward motion, with $\delta_{2}$ adjusted accordingly. Compared to the uniform parameter, cut frequency remains similar or becomes higher while velocity during cutting is reduced for carrot, courgette and potato.
These results do not provide an exhaustive characterization, which is beyond the scope of the current paper, as this would require a proper optimization framework and extensive validation tests for a proper statistical analysis, further complicated by the extreme variability in mechanical properties of vegetables, Agrawal and Lucas (2003). However, they illustrate the tunability of the virtual model controller, by showing how different parameter configurations can achieve successful cutting while producing distinct motion characteristics, allowing the control designer to shape the desired behavior through physical parameter selection rather than trajectory specification. Fig. 1 shows the sliced vegetables at the end of experiments. By visual inspection, both (i) uniform and (ii) food-specific parameter selection produce no noticeable difference in slice quality or accuracy.
@@
(a) Rocking with the proposed VMC succeeds for knives with different sizes and blade profiles. In contrast, impedance control following a pre-defined trajectory fails under variations in knife geometry. The reference trajectory is the end-effector position and orientation recorded during a successful cut with Knife 1 under VMC. Trajectories of the knife tip, heel and the end-effector are shown in each figure.
(b) Forces, velocities, and cut frequencies for different knives. Blue: cutting with the proposed VMC; red: impedance control on predefined trajectory.
Figure 10: Rocking motion with different knives using the same virtual mechanism and control parameters.
6.2 Accuracy
@@
(b) Forces and velocities for different heights of the board. Impedance control fails to complete the motion at a +24 mm height with high gains and at +32 mm with low gains due to excessive contact forces on the end-effector, while the proposed VMC continues to operate. Blue: cutting with the proposed VMC; red: impedance control (high gain); yellow: impedance control (low gain).
Figure 11: Rocking with different cutting boards using the same virtual mechanism and control parameters.
6.3 Robustness
Our virtual model controller demonstrates inherent robustness. As shown in previous section, the same set of control parameters achieved successful cut of a variety of food items, from delicate spring onions to rigid potatoes. In addition to these results, we tested the controller using two different knives. Knife 1 features an optimal blade curvature for rocking and has been used in all experiments so far. Knife 2 has a different blade geometry, which is still suitable for rocking. Results are shown in Fig. 10. To assess robustness, we consider a uniform set of parameters for both knives. Although both knives exhibited rocking motion, the differences in blade geometry affected energy dissipation and resulted in different rocking features. Knife 2 has a smaller blade than Knife 1, resulting in a reduced energy dissipation through contact with the cutting board. This led to a higher residual energy after cutting, which turned into speed, that is, faster cuts. The increase in force peaks for Knife 2 can be explained in a similar way: Knife 2 moves faster than Knife 1, leading to stronger forces around the transition from the cutting phase to the raising phase.
We compared our approach against a standard impedance controller based on pre-defined nominal trajectory. Deriving a suitable rocking trajectory is itself non-trivial, as it requires precise knowledge of the knife geometry and cutting board height. We therefore constructed the nominal trajectory from the end-effector trajectory of a successful cutting motion with Knife 1 generated by the virtual model controller. An impedance controller with high gains (Table 1 (d)) was used for tracking it, tuned manually to achieve the best accuracy without inducing high-frequency oscillations. For impedance control, we define the cutting phase as the time interval during which the end-effector $z$-position is decreasing. Cut forces are then computed using (15). With Knife 1, the rocking motion could be reproduced and cutting was generally successful, although noticeable tracking discrepancies remained. Notably, a thin film of onion at the base was left unseparated in most of the cuts, indicating insufficient normal cutting force. With Knife 2, the geometric mismatch of the knife shape caused the blade to lose contact with the board, leading to poor cuts without complete separation.
We also demonstrate the ability of the virtual model controller to handle different heights of the cutting board. Fig. 11(a) illustrates the experimental setup, where the nominal cutting board height was modified by adding MDF layers underneath. Fig. 11 shows that, without changing any control parameters, the virtual model controller maintained a stable rocking motion across all heights. As the board height increased, the peak cutting force and peak and average cut velocities gradually decreased, which can be explained by the reduced distance available for the knife to accelerate before switching and by the saturating profiles of the nonlinear springs.
Also in this scenario, we compared our approach against an impedance controller, testing high-gain and a low-gain impedance setting (Table 1 (d)). The impedance controller tracked the pre-defined nominal trajectory constructed as specified above. For the cutting board at the nominal height, the low-gain impedance controller failed to cut onions completely, with the cut approximately reaching halfway through due to insufficient applied force. When the cutting board was raised by 8 mm, the same impedance controller was then able to cut through the spring onions, as the higher board effectively increased penetration for a given trajectory. However, further increases in board height led to larger contact forces, triggering protective stops due to excessive forces on the arm. Specifically, the high-gain controller failed at 24 mm height increment, while the low-gain controller tolerated higher displacements but ultimately failed at 32 mm increment.
6.4 Portability
The proposed virtual mechanism is platform independent. We demonstrate this by implementing the same controller on a humanoid robot from RT Corp, Sciurus17. Unlike the Franka platform, Sciurus17 has a lower payload capacity, slower communication speed and uses current control as a proxy of torque control. These differences can be managed by re-tuning the parameters of the controller to ensure smaller forces and slower motions. The cutting plane was kept fixed, and food items were manually fed to the knife to reduce system burden. Fig. 12 shows that the inherent limitations of the robotic platform led to reduced performance. However, successful rocking motion is achieved.
@@ 6.4 Portability
7 Conclusion and future works
We have proposed a new virtual model controller for rock cutting. The rocking motion has been discussed theoretically, through energy considerations, and experimentally. Performance has been illustrated by cutting several vegetables, for several requirements of thicknesses and cut frequencies. Robustness has been illustrated by using two different knives and several cutting board heights. Figs. 10 and 11 clarify the difference between the virtual model controller and other control methods based on tracking of a nominal, preplanned trajectory. The latter typically require precise knowledge of knife geometry and high-gains for accurate tracking, which limit robustness. In contrast, the rocking motion established by the virtual model controller is not pre-programmed but shaped by the interaction between robot, controller and environment, therefore robustly adapts to a wide range of settings. The experimental evaluation of the impedance controller show the limits of shaping the end-effector impedance without explicitly taking into account the geometry of the task. In contrast, the design of the virtual model controller starts from the task geometry to structure the compliance of the virtual mechanisms in a task-relevant way. This leads to more flexible tuning and increased robustness. For instance, the maximum force $\sigma_{1}$ affects directly the contact force with the cutting board. Likewise, the maximum force $\sigma_{\mathrm{ori}}$ penalizes deviations from the slicing plane. Experiment videos are available in Extension 1.
Successful operation depends on reasonable selection of parameters for the virtual model controller. The position of the object and the orientation of the knife relative to the end effector must be known to ensure effective cutting. However, precise geometric details of the knife or the exact height of the cutting board are not required. Similarly, the exact attachment points of the virtual components are not critical, as approximate placement is sufficient. These features highlight critical differences with Mu et al. (2023), which is the only prior work explicitly considering rolling motion, to the best of our knowledge. In Mu et al. (2023), the motion is decomposed into three stages (pressing, touching, and slicing) combined to different control modes (position, hybrid position–impedance and hybrid position–force control). The implementation requires switching among multiple controllers to track motion and handle contact at different stages of the cut, and fundamentally relies on the explicit geometric model of the knife to plan contact-point transitions and compute associated Jacobians. Mu et al. (2023) demonstrates single-cut performance on selected food items but does not consider repetitive cutting cycles, so robustness under repeated operation and environment variations is not evaluated. In contrast, our virtual-model controller manages both motion generation and interaction without mode switching and operates with substantially less information.
The proposed design is a key component in the development of a fully integrated cutting pipeline. This involves autonomous pick and place of the food, its orientation on the cutting board, and the ability to hold the food in place during cutting. The robot should also be able to recognize different foods and knives, and to measure in real-time the result of its action, during and after cutting. At the level of food-knife interaction, a key future direction is the systematic exploration of adaptation/learning of the control parameters, for stronger performance. The geometric configuration of the virtual model controller serves as the primary driver of the robot’s behavior, while parameters such as stiffness and damping are used for fine-tuning. This provides a fundamental advantage for tuning, as the geometry defines the core motion while the control parameters refine it. The objective is to exploit this structural advantage in combination with external sensing, such as vision and tactile feedback, to enable real-time parameter adaptation. For example, the stiffness $k_{2}$ can be adjusted in real-time to achieve successful food separation while moderating the force exerted by the knife on food and board. The short experimental study proposed below show encouraging preliminary results. Every six cutting cycles we increase $k_{2}$ if the knife does not achieve separation of food and we reduce $k_{2}$ if the knife’s handle overshoot is too large (i.e. if the knife handle moves past the critical reference). This leads to the adaptation law
@@ 7 Conclusion and future works
where $e$ is the error signal capturing the mismatch between desired and achieved vertical position of the knife, and $\alpha$ is the adaptation gain. Results are summarized in Fig. 13. $k_{2}$ starts from a low value, which is sufficient to cut through a carrot. From time $t=20$s, the carrot presents a harder consistency which requires adaptation. Successful cutting is restored after time $t=35s$ through adaptation of $k_{2}$.
@@
Table 1: Control parameters for each experiment. Units: stiffness $k$ [N/m], maximum force $\sigma$ [N], damping $c$ [N·s/m], position $\boldsymbol{r}$ [m], mass $m$ [kg], $\delta$ [m], $\tau$ [s]. The mass–spring–damper system ($m_{b},k_{b},c_{b}$) generates smooth transitions between reference points. Parameters are selected to ensure overdamped behavior, and the time constant $\tau$ of the slower pole is reported as an intuitive measure of the transition dynamics. Consistent parameters across all experiments: masses for simulating the sliding motion ($m_{a}=m_{\text{ori}}=0.1$). Consistent parameters across all experiments on Franka: $\sigma_{1}=10,c_{1}=1,\sigma_{\text{ori}}=50,c_{\text{ori}}=10$, and $\boldsymbol{r}_{2,1}\,(y,z)=0.43,-0.1$.
(a) Uniform control parameters and parameters tuned to reduce force for individual vegetables. The most sensitive parameters here are the stiffness values ($k_{1}$, $k_{2}$) governing contact and cutting. $m_{b},k_{b},c_{b}$ regulate the transition between reference points $\boldsymbol{r}_{2,1}$ and $\boldsymbol{r}_{2,2}$ to maintain similar cutting frequency when $k_{2}$ is reduced. $\boldsymbol{r}_{2,2}$ and $\delta_{2}$ are adjusted according to object height. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Uniform parameters 25 150 25 2 1200 0.58, 0.4 0.02 0.15 0.5
Spring onion 10 65 25 2 1200 0.58, 0.4 0.02 0.2 0.05
Cucumber 10 75 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Courgette 10 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Carrot 20 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
Potato 20 105 25 2 1200 0.83, 0.5 0.02 0.25 0.05
(b) Different cuttiing frequencies. The primary tuning parameter is $k_{2}$, increased for faster cuts. $m_{b}$, $k_{b}$, and $c_{b}$ are adjusted to switch faster, while $k_{1}$ is increased to ensure stable contact during rapid motion. $\delta_{1}$ is increased for faster cuts to accommodate higher knife velocity and avoid overshoot during motion reversal. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Fast cut (0.94 cut/s) 45 200 30 2 800 0.73, 0.35 0.035 0.145 0.015
Slow cut (0.51 cut/s) 25 120 25 2 800 0.58, 0.4 0.02 0.15 0.5
(c) Robustness to environment variations. Identical mechanism and control parameters are used across all robustness experiments. Robot: Franka.
Exp $k_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
Robustness 30 130 25 2 800 0.63, 0.4 0.01 0.2 0.1
(d) Robustness to environment variations. Impedance gains used for the trajectory-tracking baseline. Translational stiffness is given along the Cartesian axes, and rotational stiffness about the corresponding axes. Robot: Franka.
Translational stiffness Rotational stiffness
$K_{x}$ $K_{y}$ $K_{z}$ $K_{r_{x}}$ $K_{r_{y}}$ $K_{r_{z}}$
High gain 1500 7000 7000 300 170 170
Low gain 750 3500 3500 150 85 85
(e) Platform independence. Parameters are adjusted to account for differences in actuation capability and communication frequency between platforms. Stiffness and max forces are reduced. Robot: Sciurus.
$k_{1}$ $\sigma_{1}$ $c_{1}$ $k_{2}$ $\sigma_{2}$ $c_{2}$ $k_{\text{ori}}$ $\sigma_{\text{ori}}$ $c_{\text{ori}}$ $\boldsymbol{r}_{2,1}\,(y,z)$ $\boldsymbol{r}_{2,2}\,(y,z)$ $\delta_{1}$ $\delta_{2}$ $\tau$
10 2 0.3 20 10 0.1 30 10 1 0.4, -0.1 0.6, 0.6 0.047 0.38 0.45
{acks}
Yi Zhang is supported by the Engineering and Physical Sciences Research Council, AgriFoRwArdS Center for Doctoral Training [EP/S023917/1] and RT Corp.